OpenStudy (anonymous):

Simplify 68 – 18 ÷ 3 • 6

OpenStudy (xishem):

Order of operations: 1) Parentheses 2) Exponents 3) Mult/Div 4) Add/Sub Using that:\[68-18/3*6=68-((18/3)*6)=68-((6)*6)=68-(36)=32\]